OpenStudy (anonymous):

CAN ANYONEEEEEEEEEEE integrate : Xlogsinx dx from 0 to pi

OpenStudy (anonymous):

no it doesnot show step by step

OpenStudy (dumbcow):

hmm indefinite integral can't be done using elementary functions

hartnn (hartnn):

use the property \(\large \int \limits_a^bf(x)dx=\int \limits_a^bf(a+b-x)dx\) and see how much it simplifies!

hartnn (hartnn):

ofcourse you have to add the 2 equations.....

hartnn (hartnn):

I =int Xlogsinx dx --------->(1) I =int (pi- X)logsinx dx ----->(2) ADD (1) and (2) what u get ?

OpenStudy (anonymous):

pi logsinx

OpenStudy (anonymous):

2i=pilogsinx dx

hartnn (hartnn):

yes, now you can refer this : http://openstudy.com/users/aliza_k#/updates/5039a6e2e4b043c156a31ff9

OpenStudy (dumbcow):

?? im lost \[\int\limits_{0}^{\pi} x \ln(\sin x) dx \ne \int\limits_{0}^{\pi} \pi \ln(\sin x) dx\]

hartnn (hartnn):

its 2I = int pi*log sin

OpenStudy (dumbcow):

oh ok

hartnn (hartnn):

doubts ? i have used that property ...and then added....ask if you still didn't get it...

OpenStudy (anonymous):

no doubt thank u so much @hartnn

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

omg it's too lengthy

hartnn (hartnn):

yes, it is. What you have posted is not an easy integral to solve. But after using that property, it reduces to classic definite integral problem (ln sin x), which has a classic solution i explained there.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

my answer is cmg -pi/2 (log2)

hartnn (hartnn):

there, I =I/2 -pi/2 ln 2 so, I/2 = pi/2 ln 2 I = pi ln 2

OpenStudy (anonymous):

-pi^2/2 (log2)

hartnn (hartnn):

oh, wait.... yes thats correct -pi^2/2 (log2) = -3.4205 which wolf gave....